OpenStudy (anonymous):

Can someone show me how to do this please ? Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 2x – 5y = –12 3x – 4y = 3

OpenStudy (anonymous):

Equation 1 : 2x -5y = -12 Equation 2 : 3x - 4y = 3 Multiply Equation 1 by (3) : 6x - 15y = -36 Multiply Equation 2 by (2) : 6x - 8y = 6 can you carry on :D ?

OpenStudy (anonymous):

Why do u add them if your trying to eliminate ?

OpenStudy (anonymous):

I am making the " x " the same so I can eliminate them

OpenStudy (anonymous):